INAF queue observing, 2015 Mar 16-17

Observer: R. Speziali, S. Bisogni
Telescope Operator: G. Bechetti
Support Astronomer: Olga Kuhn
Instrument: MODS/LBC

Summary

At the sunset the sky is clear but humidity is too high to open.

Night duration (12\xB0 el twilight) 19:30-05:30 = 11 hrs

Observing time: 7 hrs

Weather time loss: 4 hrs (37%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
SARACCO C RDCS mask 2 red grating 1h average seeing 1.4''
Standard feige34 red and dual grating 12min average seeing 1.4''
FIORE B GOODS-North Field 2 dual grating 1h15min average seeing 1.4'', 45 min out of spec for diff refraction?
NASCIMBENI WASP-39 Parallel 2h 54/500 spectra in red, 34/500 in blue, 1h with good sky, then clouds
MODS Calibrations

Details

02:50 TO opens the dome and starts collimating

- SARACCO C -

03:00 start acquisition of Saracco 19C = mods1r.20150317.0016-0018=

03:22 execMODS mods.11.RDCSm2_1.obs mods1r.20150317.0019-0022
  • Seeing = 1.3"-1.5"
  • seeing improved between second an third integrations = 1.2''
  • during the third integration seeing get worse again = 1.4''-1.6''
  • during the last integration seeing is getting better = 1.2''-1.3''

04:30 start acquisition of FEIGE34 (red & dual grating) mods1r.20150317.0023-0024

04:43 execMODS FEIGE34.obs mods1r.20150317.0025-30, mods1b.20150317.0015-18

04:50 dofpia (parallel) on the red channel

05:00 the seeing is dropping to 1''

- FIORE B -

05:11 slewing to FioreB

05:13 acqMODS goodsnorth2_mos.acq mods1r.20150317.0031-33
  • seeing 1.5''

05:21 dofpia, /x2, /redonly

05:25 execMODS goodsnorth2_mos_1.obs mods1r.20150317.0034-37, mods1b.20150317.0019-22
  • seeing 1.3''

05:27 Launched LBC red OB in parallel mode

06:35 execMODS goodsnorth2_mos_2.obs (15min ONLY!) mods1r.20150317.0038, mods1b.20150317.0023
  • seeing 1.3''

06:35 dofpia, /x2, /redonly

06:42 Launched LBC red OB in parallel mode

- NASCIMBENI -

06:53 Acquisition of WASP39 field of Nascimbeni mods1r.20150317.0039-
  • seeing 1.2'

06:53 Acquisition stopped: Rotator Unwrap

06:58 Acquisition of WASP39 field of Nascimbeni mods1r.20150317.0040-41
  • seeing 1.2'

07:13 execMODS mods.wasp39.obs mods1r.20150317.0042-4
  • seeing 1.2'
  • seeing 1.6'' during the 10th exposition

07.27 launched wasp39_lbc.ob
  • wasp39 is saturated, we defocus more
  • seeing 1.6''
  • coma problems with LBC, TO tries corrections
  • run dofpia again
  • 07:46 seeing 1.7''
  • not able to reach the requested counts, so we change the exposure time (t_new=t_old*0.60)
  • 08:47 seeing = 1.4''
09.15 Clouds are passing. MODS Guiding and LBC images are gone. We keep on with MODS.

09.25 We stopped acquiring with LBC.

10.05 We stopped acquiring with MODS.

10.10 Resuming acquisition with MODS.

10:45 Pausing the acquisition

11:00 resuming again the acquisition

11:07 Pausing again the acquisition, no spectra in the slits

11:49 it's too cloudy. The TO closes the dome so we can run MODS calibrations.

12:02 We run MODS calibrations
  • execMODS NGC4535_mods.cal Sommariva mods1r.20150317.0094-98 == ==mods1b.20150317.0059-63
  • execMODS mods.11.RDCSm2.cal SaraccoC mods1r.20150317.0099-0102
  • execMODS goodsnorth1_mos_arcflat.cal FIORE_A mods1r.20150317.0103-0107 mods1b.20150317.0064-0068
  • execMODS goodsnorth2_mos_arcflat.cal FIORE_B mods1r.20150317.0108-112 mods1b.20150317.0069-0073
  • execMODS mods.wasp39.cal Nascimbeni mods1r.20150317.0113-0117 mods1b.20150317.0074-0078
  • execMODS mods.20..cal Mignoli MOS20 mods1r.20150317.0118-0121

-- %USERSIG{SusannaBisogni - 2015-03-16}%

Comments


 
Topic revision: r18 - 29 Apr 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback